Macy wants to change 18 quarts of an 8% solution of ammonia to a 6% solution by adding pure water. How much pure water must she add?
Answer by stanbon(75887)

Macy wants to change 18 quarts of an 8% solution of ammonia to a 6% solution by adding pure water. How much pure water must she add?
------------------
Equation:
water + water = water
-----
0.92*18 + 1*x = 0.94(18+x)
--------
92*18 + 100x = 94*18 + 94x
6x = 2*18
x = 6 qts (amt. of pure water to add)
------------------
Cheers,
Stan H.
--------------------